solve the next ecuation with laplace transformation (find y(t))

u(e– cm = [etyle – 7) dt.
Answer #1

+ Ylt) - ezt et y (t-1)dt >> y(t) - ezt = etry (U) ( By Convolution formula) toy = Taking Zaplace trunsform, both sides L{y(tTaking Inverse zaplace transform both sides, we get w =) LY{Y(8) 2* { (5-2) 7 - S = (5=23 (5-2) (18-27 ezt t + ezt - 2t it e